OpenStudy (anonymous):

In October, Greg and Thomas had the same amount of money in their savings accounts. In November, Greg deposited $120 into his account. Thomas increased the money in his account by 20%. When they compared their balances, they found that they were still equal. How much money did each person have in his account in October?

OpenStudy (anonymous):

@ccastudentgrade10

OpenStudy (anonymous):

@rugerg

OpenStudy (anonymous):

@rugerg

OpenStudy (anonymous):

@ccastudentgrade10

OpenStudy (anonymous):

If Greg deposited $120 into his account and Thomas increased his money by 20% and they still got the same amount in November, it means that $120 is 20% of their money in October. 120 = 20% * x (where x is the money in October) 120 = 20/100x 120 = 2/10x 120 = 1/5x / * 5 (both sides) x = 600 So each of them had $600 in October. Doublecheck: Greg's money = 600 + 120 = 720 Thomas' money = 600 + 20% * 600 = 600 + 120 = 720

OpenStudy (anonymous):

omj thx ur the best